# Learning out-of-time-ordered correlators with classical kernel methods

### v1

@@ III Methods
The manner in which information propagates through the system $\mathcal{S}$ depends on the details of its dynamics, which is specified by the time-independent Hamiltonian $H$. This means that in order to compute OTOCs and related quantities, such as the sum of OTOCs in (17), generally we need to simulate the evolution of the system $\mathcal{S}$. This amounts to calculating $e^{-iHt}$. Such a calculation becomes computationally intensive for a generic $H$ as the size of the system increases, and needs to be performed for every $H$ and value of time $t$ that one wishes to investigate. In spite of this, is there some property of many-body Hamiltonians that can be used to determine OTOCs without the need to simulate the full dynamics? Here we investigate whether such properties may be captured by a ML model which makes use of KMs as described in Algorithms 1 and 2. In this section we describe the methods used to conduct the investigation. This includes a description of the parameterised sets of Hamiltonians we consider, how the task is framed as a learning problem, the specific kernels which are trained and applied to the data, and the techniques used to generate the data.
@@ III.1 The parameterised sets of Hamiltonians
The learning problem addressed in this work requires us to choose a parameterised set of Hamiltonians $\{H(x):x\in\mathbb{R}^{d}\}$ to investigate. This determines both the set of possible dynamics for $\mathcal{S}$, and the explicit form of the OTOCs considered in the learning problem. With this in mind, we consider 4 distinct choices of the parameterised set of $n$-qubit Hamiltonians. Each of the sets forms a 3-dimensional (i.e., $d=3$) subspace of Hermitian operators describing a collection of 1D quantum systems. The parameterised sets are chosen to contain systems which are widely considered in condensed matter physics.
@@
for all $x=(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}$. Similar to $\mathscr{H}_{2}$, the Hamiltonians in $\mathscr{H}_{3}$ provide quantum models of ferromagnetic and anti-ferromagnetic materials. They also describe 1D spin-chains with nearest-neighbour spin couplings, in addition to next-nearest neighbour spin-couplings, where the relative magnitude of the next-nearest-neighbour couplings are half as great as the nearest-neighbour couplings. This is a slightly more realistic model of anti-ferromagnetic and ferromagnetic materials which does not assume that atoms only interact with their nearest-neighbours.
@@
for all $x=(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}$. The Hamiltonians in this set describe a semi-classical version of the 1D Ising chain under the influence of an external homogenous magnetic field with non-zero components in the physical $x$ and $z$ directions.
With reference to equations (18), (19), (20) and (21), when we evolve a quantum system according to the unitary operator $e^{-iH_{j}(x)}$ (where $j\in\{1,2,3,4\}$), the evolution time $t$ is effectively quantified by the Euclidean norm $||x||$ of the parameter vector $x\in\mathbb{R}^{3}$. This means that limiting the evolution time of the associated quantum systems corresponds to restricting the norm of the parameter vector $x\in\mathbb{R}^{3}$. Accordingly, it is most natural to sample the input data (i.e., the parameter vectors) from a ball in $\mathbb{R}^{3}$ centered at the origin, whose radius determines the maximum evolution time.
@@
For $\mathscr{H}_{1}$, $\mathscr{H}_{2}$ and $\mathscr{H}_{3}$, we chose to uniformly sample input data from the ball of radius $n$ centered at the origin. This means that the effective evolution time $t$ of the corresponding $n$-qubit quantum systems is bounded above by $t\leq n$. This is motivated by both computational budgets and empirical observations from initial simulations, which showed that linear time in the system size is sufficient to exhibit interesting behaviours or dynamics in the system (see Figures 4(a)–(c)). For $\mathscr{H}_{4}$, we chose to uniformly sample input data from the ball of radius $2n$ centered at the origin. Increasing the radius of the ball for $\mathscr{H}_{4}$ was motivated by the observation that sampling from the ball of radius $n$ did not produce much variation in the value of the OTOCs we consider (see Figure 4(d)). Informally, one possible explanation for this is that the only term in (21) which contributes to the spread of information is the term multiplied by $x_{3}$, and this term only transfers information about $Z$ to neighbouring qubits. In contrast, the terms in the Hamiltonians in $\mathscr{H}_{1}$, $\mathscr{H}_{2}$ and $\mathscr{H}_{3}$ that contribute to the spread of information collectively transfer at least $Z$ and $X$ information. Accordingly, it seems reasonable that the systems in $\mathscr{H}_{4}$ require a greater maximum evolution time to observe changes in the correlations captured by the OTOC.
[…]
Figure 4: Scatter plot of the 40-qubit training data for $\mathscr{O}_{Sum}$. The colour of a point at $(x_{1},x_{2},x_{3})$ specifies the value of $\mathscr{O}_{Sum}(x_{1},x_{2},x_{3})$ as calculated with the MPO-based algorithm. The plots display the $\mathscr{O}_{Sum}$ training data for $\mathscr{H}_{1}$ in (a), $\mathscr{H}_{2}$ in (b), $\mathscr{H}_{3}$ in (c), and $\mathscr{H}_{4}$ in (d).
@@ III.2 The learning problem
for all $x\in\mathbb{R}^{d}$, where we denote by $\mathcal{Z}_{1}(x)\equiv e^{iH(x)}Z_{1}e^{-iH(x)}$ the operator $Z_{1}$ evolved in the Heisenberg picture under $H(x)$ for one unit of time. The Hamiltonian parametrisations that we consider have the property $H(tx)=tH(x)$ for any $t\in\mathbb{R}$, hence we simply absorb the evolution time $t$ into the Euclidean norm of $x\in\mathbb{R}^{d}$.
The second function, denoted $\mathscr{O}_{Sum}:\mathbb{R}^{d}\to\mathbb{R}$, captures the sum of OTOCs in (17) when $q_{A}$ and $q_{B}$ lie at opposite ends of the 1D spin chain, and is given by
@@
The learning task is then to find a model which accurately learns $\mathscr{O}_{XZ}$ or $\mathscr{O}_{Sum}$ from sampled data (we describe our procedure for generating data in Section III.4). In total, we consider 64 instances of the learning task: one for each choice of $\mathscr{O}_{XZ}$ or $\mathscr{O}_{Sum}$, with one of the four parameterised sets of Hamiltonians $\mathscr{H}$ from Section III.1, and a system size of 5, 10, 15, 20, 25, 30, 35 or 40 qubits. For each of the tasks, we examine 6 different choices of kernel functions (see Section III.3) and compare their performance by determining how well the models perform on unseen test data. Thus, the generated data is split into training and testing data, 80% for training and 20% for testing. Next we perform a 10-fold cross validation on the training data to find suitable hyperparameter values for the various kernels using a grid-search. The hyperparameter values which we trial are listed in Table I of Appendix B. The best hyperparameter values are then selected based on the average coefficient of determination ($R^{2}$) over all 10 folds of the training data, and are used to train and apply ML models to the testing data. The best hyperparameter values for every instance of the learning task can be found in Tables IIV of Appendix B.
@@ III.3 The kernels
We now provide definitions of the 6 kernels which are used to train ML models and make predictions on the testing sets for each problem instance. Some of the kernels depend on hyperparameters $\gamma\in\mathbb{R}$, $c_{0}\in\mathbb{R}$ and $d\in\mathbb{N}$ which, as discussed above, are tuned via a 10-fold cross-validation on the training data (see Section III.2). The first kernel is the linear kernel, denoted $\mathcal{K}_{\textrm{lin}}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$, which is defined by the standard Euclidean inner product of its arguments,
@@
Each of these kernels are commonly used in ML algorithms that apply KMs to regression problems, which lead to our use of them here. Additionally, the Scikit Learn Python package [74] has a function pairwise_kernels for conveniently calculating the associated kernel matrices. Using this function, the 10-fold cross-validation for the regularisation strength $\lambda$ and the other hyperparameters $\gamma\in\mathbb{R}$, $c_{0}\in\mathbb{R}$ and $d\in\mathbb{N}$ is performed.
@@ III.4 Generating the data
To generate the data for this work, we use an efficient numerical algorithm based on matrix product operators (MPOs) [41]. Specifically, we sample inputs $x$ from subsets of $\mathbb{R}^{3}$ (see Section III.1), which effectively selects a Hamiltonian $H(x)$ from one of the sets $\mathscr{H}$. We then directly calculate the associated labels determined by $\mathscr{O}_{XZ}$ and $\mathscr{O}_{Sum}$ with the MPO-based algorithm. This involves representing each of the operators under the trace in (22) and (23) as an MPO, and evolving them in the Heisenberg picture using time-evolving block decimation (TEBD) [75, 76].
@@
Applying the algorithm requires us to choose an appropriate maximum bond dimension $\chi$ for the MPOs. A larger $\chi$ makes the algorithm more computationally expensive to implement, but also means that the computed numerical values are generally more accurate. Accordingly, we need to pick a bond dimension that both fits within the computational budget, and leads to reasonably accurate labels for the data points. In Figure 5, we provide plots of some of the labels for the 40-qubit systems against the value of $\chi$ used to calculate them, which shows how well the labels have converged with respect to the bond dimensions selected for producing the datasets. In an attempt to balance the trade-off between compute costs and convergence, we chose to use a maximum bond dimension $\chi$ of 110, 140, 110 and 150 to calculate the values of the function $\mathscr{O}_{XZ}$ with $\mathscr{H}_{1}$, $\mathscr{H}_{2}$, $\mathscr{H}_{3}$ and $\mathscr{H}_{4}$, respectively. Similarly, we chose to use a maximum bond dimension $\chi$ of 70, 90, 70 and 100 to calculate the values of the function $\mathscr{O}_{Sum}$ with $\mathscr{H}_{1}$, $\mathscr{H}_{2}$, $\mathscr{H}_{3}$ and $\mathscr{H}_{4}$, respectively.
[…]
Figure 5: The numerical values of $\mathscr{O}_{XZ}$ and $\mathscr{O}_{Sum}$ for 10 randomly sampled inputs plotted against the maximum bond dimension $\chi$ used to calculate them for 40-qubit quantum systems. Plots (a), (b), (c) and (d) show $\mathscr{O}_{XZ}$ with $\mathscr{H}_{1}$, $\mathscr{H}_{2}$, $\mathscr{H}_{3}$ and $\mathscr{H}_{4}$, respectively. Plots (e), (f), (g) and (h) show $\mathscr{O}_{Sum}$ with $\mathscr{H}_{1}$, $\mathscr{H}_{2}$, $\mathscr{H}_{3}$ and $\mathscr{H}_{4}$, respectively. The vertical black lines indicate the value of $\chi$ used to calculate the rest of the data for the corresponding function and parameterised set of Hamiltonians. In the same order as the plots, the chosen values of $\chi$ are 110, 140, 110, 150, 70, 90, 70 and 100.
@@
In addition to this selection of bond dimensions, we use a Trotter step size of $\Delta t=0.05$ to perform the Heisenberg picture time evolution of the MPOs using TEBD. This seems reasonable when one considers that the magnitude of the error terms involved in the algorithm scale as $\mathcal{O}\left((\Delta t)^{3}\right)$. We also make use of a time-splitting method (see Section IV.B.2 in [15]) to improve the numerical accuracy of the algorithm by reducing the amount of time that the MPOs need to be evolved for. Then for each qubit number considered (5, 10, 15, 20, 25, 30, 35 and 40), each parameterised sets of Hamiltonians, and both functions $\mathscr{O}_{XZ}$ and $\mathscr{O}_{Sum}$, we generate 1000 data samples for training and 250 for testing. The random samples of the inputs are drawn using the same random seed. Thus, for each qubit number, all datasets contains the same input parameters. The values of the functions are then computed using the ITensor Julia library [77] to implement the MPO-based time-splitting algorithm. It should be noted though, that even with the MPO-based algorithm, producing the data in this work required $\sim 10^{6}$ CPU hours, and thus required access to a supercomputer.

### v2

@@ III Methods
The manner in which information propagates through the system $\mathcal{S}$ depends on the details of its dynamics, which is specified by the time-independent Hamiltonian $H$. This means that in order to compute OTOCs and related quantities, such as the sum of OTOCs in (18), generally we need to simulate the evolution of the system $\mathcal{S}$. This amounts to calculating $e^{-iHt}$. Such a calculation becomes computationally intensive for a generic $H$ as the size of the system increases, and needs to be performed for every $H$ and value of time $t$ that one wishes to investigate. In spite of this, is there some property of many-body Hamiltonians that can be used to determine OTOCs without the need to simulate the full dynamics? Here we investigate whether such properties may be captured by a ML model which makes use of KMs. In this section we describe the methods used to conduct the investigation. This includes a description of the parameterised sets of Hamiltonians we consider, how the task is framed as a learning problem, the specific kernels which are trained and applied to the data, and the techniques used to generate the data.
@@ III.1 The parameterised sets of Hamiltonians
The learning problem addressed in this work requires us to choose a parameterised set of Hamiltonians $\{H(x):x\in\mathbb{R}^{d}\}$ to investigate. This determines both the set of possible dynamics for $\mathcal{S}$, and the explicit form of the OTOCs considered in the learning problem. With this in mind, we consider four distinct choices of the parameterised set of $n$-qubit Hamiltonians. Each of the sets forms a 3-dimensional (i.e., $d=3$) subspace of Hermitian operators describing a collection of 1D quantum systems. The parameterised sets are chosen to contain systems which are widely considered in condensed matter physics.
@@
for all $x=(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}$. Similar to $\mathscr{H}_{2}$, the Hamiltonians in $\mathscr{H}_{3}$ provide quantum models of ferromagnetic and anti-ferromagnetic materials. They also describe 1D spin-chains with nearest-neighbour spin couplings, in addition to next-nearest neighbour spin-couplings with half the coupling strength. This is a slightly more realistic model of anti-ferromagnetic and ferromagnetic materials which does not assume that atoms only interact with their nearest-neighbours.
@@
for all $x=(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}$. The Hamiltonians in this set describe 1D spin-chains with nearest-neighbour $Z$-component spin couplings under the influence of an external homogenous magnetic field with non-zero components in the physical $x$ and $z$ directions.
The parameterised Hamiltonians we consider all satisfy the property $H(tx)=tH(x)$ for any $t\in\mathbb{R}$, where $H$ denotes one of the parameterised Hamiltonians defined in equations (19)–(22). Using this property, but replacing $t\mapsto\|x\|$ and $x\mapsto\hat{x}$ (where $\hat{x}=x/\|x\|$ is the unit vector parallel to $x$), we see that $H(x)=\|x\|H(\hat{x})$, which further implies that
$e^{-iH(x)}=e^{-i\|x\|H(\hat{x})}.$ (23)
Equation (23) shows how the operator $e^{-iH(x)}$ is equivalent to the time-evolution operator describing evolution under the Hamiltonian $H(\hat{x})$ for a time $t=\|x\|$. This holds for all parameterised Hamiltonians defined in (19)–(22), and means that we can capture changes in time by scaling the parameter vector $x$, which changes $\|x\|$ but leaves $H\left(\hat{x}\right)$ fixed (up to a sign if we scale by a negative constant). Hence, choosing a range of evolution times for the associated quantum systems corresponds to choosing the parameter vectors with norms lying in the same range.
Accordingly, it is most natural to sample the input data (i.e., the parameter vectors $x$) from a ball in $\mathbb{R}^{3}$ centered at the origin, whose radius determines the maximum evolution time. The goal is then to choose an appropriate radius for each of the balls which is both feasible to simulate and captures an interesting range of evolution times. Specifically, we chose the radius of the ball from which the input data is sampled to scale linearly with the size of the underlying quantum system. As discussed in Section II.3, the scrambling time for qubit models scales linearly in the size of the system, so we expect that a linear scaling (in the system size) of the maximum evolution time will be sufficient to capture both the short-time regime and possibly part of the intermediate-time regime.
@@
For $\mathscr{H}_{1}$, $\mathscr{H}_{2}$ and $\mathscr{H}_{3}$, we chose to uniformly sample input data from the ball of radius $n$ centered at the origin. This means that the effective evolution time $t$ of the corresponding $n$-qubit quantum systems is bounded above by the system size $t\leq t_{max}=n$. For $\mathscr{H}_{4}$, we chose to uniformly sample input data from the ball of radius $2n$ centered at the origin, meaning that $t\leq t_{max}=2n$. Increasing the radius of the ball for $\mathscr{H}_{4}$ was motivated primarily by the observation that sampling from the ball of radius $n$ did not produce much variation in the value of the OTOCs. Informally, one possible explanation for this is that the only term in (22) which contributes to the spread of information is the term multiplied by $x_{3}$, and this term only transfers information about $Z$ to neighbouring qubits. In contrast, the terms in the Hamiltonians in $\mathscr{H}_{1}$, $\mathscr{H}_{2}$ and $\mathscr{H}_{3}$ that contribute to the spread of information collectively transfer at least $Z$ and $X$ information. Accordingly, it seems reasonable that the systems in $\mathscr{H}_{4}$ require a greater maximum evolution time to observe changes in the correlations captured by the OTOC.
@@ III.2 The learning problem
for all $x\in\mathbb{R}^{d}$, where we denote by $\mathcal{Z}_{1}(x)\equiv e^{iH(x)}Z_{1}e^{-iH(x)}$ the operator $Z_{1}$ evolved in the Heisenberg picture under $H(\hat{x})$ for $\|x\|$ units of time (see Section III.1). The second function, denoted $\mathscr{O}_{Sum}:\mathbb{R}^{d}\to\mathbb{R}$, captures the sum of OTOCs in (18) when $q_{A}$ and $q_{B}$ lie at opposite ends of the 1D spin chain, and is given by
@@
The learning task is then to find a model which accurately learns $\mathscr{O}_{XZ}$ or $\mathscr{O}_{Sum}$ from sampled data (we describe our procedure for generating data in Section III.4). In total, we consider 64 instances of the learning task: one for each choice of $\mathscr{O}_{XZ}$ or $\mathscr{O}_{Sum}$, with one of the four parameterised sets of Hamiltonians $\mathscr{H}$ from Section III.1, and a system size of $n\in\{5,10,15,20,25,30,35,40\}$ qubits. For each of the tasks, we trial six different kernel functions (see Section III.3) and compare their performance by determining how well the models perform on unseen test data. Thus, the generated data is split into training and testing data, 80% for training and 20% for testing. Next we perform a 10-fold cross-validation (see Section 4.5 of [54]) on the training data to find suitable hyperparameter values for the various kernels using a grid search. The hyperparameter values which we trial are listed in Table C.1 of Appendix C. Of the trialled values, the best value of each hyperparameter is selected based on the average coefficient of determination ($R^{2}$) over all ten folds of the training data. These hyperparameter values are then used to train ML models on the complete training dataset, which are subsequently applied to the testing data. The best trialled hyperparameter values for every instance of the learning task can be found in Tables C.2, C.3, C.4, and C.5 of Appendix C.
@@ III.3 The kernels
We now provide definitions of the six kernels which are used to train ML models and make predictions on the testing sets for each problem instance. Some of the kernels depend on hyperparameters $\gamma\in\mathbb{R}$, $c_{0}\in\mathbb{R}$ and $d\in\mathbb{N}$ which are tuned via a 10-fold cross-validation on the training data (see Section III.2).
The first kernel is the linear kernel, denoted $\mathcal{K}_{\textrm{lin}}:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$, which is defined by the standard Euclidean inner product of its arguments,
@@
Each of these kernels are commonly used in ML algorithms that apply KMs to regression problems, which lead to our use of them here. Additionally, the Scikit Learn Python package [75] has a function pairwise_kernels for conveniently calculating the associated kernel matrices. Using this function, the 10-fold cross-validation for the regularisation strength $\lambda$ and other hyperparameters $\gamma$, $c_{0}$ and $d$ is performed.
@@ III.4 Generating the data
To generate the data for this work (see Data Availability section to access the data), we use an efficient numerical algorithm based on matrix product operators (MPOs) [42], which are classical tensor networks well suited to simulating local 1D quantum systems. Specifically, we sample inputs $x$ from subsets of $\mathbb{R}^{3}$ (see Section III.1), which effectively selects a Hamiltonian $H(x)$ from one of the sets $\mathscr{H}$. We then directly calculate the associated labels determined by $\mathscr{O}_{XZ}$ and $\mathscr{O}_{Sum}$ with the MPO-based algorithm. This involves representing each of the operators under the trace in (24) and (25) as an MPO, evolving them in the Heisenberg picture using time-evolving block decimation (TEBD) [76, 77], then contracting the MPOs together to calculate the OTOC.
@@
Applying the algorithm requires us to choose an appropriate maximum bond dimension $\chi$ for the MPOs. A larger $\chi$ makes the algorithm more computationally expensive to implement, but also means that the computed numerical values are generally more accurate. Accordingly, we need to pick a bond dimension that both fits within the computational budget, and leads to reasonably accurate labels for the data points. In Figure B.1 of Appendix B, we provide plots of some of the labels for the 40-qubit systems against the value of $\chi$ used to calculate them. This shows how well the labels have converged with respect to the bond dimensions selected for producing the datasets. In an attempt to balance the trade-off between compute costs and convergence, we chose to use a maximum bond dimension $\chi$ of 110, 140, 110 and 150 to calculate the values of the function $\mathscr{O}_{XZ}$ with $\mathscr{H}_{1}$, $\mathscr{H}_{2}$, $\mathscr{H}_{3}$ and $\mathscr{H}_{4}$, respectively. Similarly, we chose to use a maximum bond dimension $\chi$ of 70, 90, 70 and 100 to calculate the values of the function $\mathscr{O}_{Sum}$ with $\mathscr{H}_{1}$, $\mathscr{H}_{2}$, $\mathscr{H}_{3}$ and $\mathscr{H}_{4}$, respectively.
@@
In addition to this selection of bond dimensions, we use a Trotter step size of $\Delta t=0.05$ to perform the Heisenberg picture time evolution of the MPOs using TEBD. This seems reasonable when one considers that the magnitude of the error terms involved in the algorithm scale as $\mathcal{O}\left((\Delta t)^{3}\right)$. We also make use of a time-splitting method (see Section VI.B.2 in [15]) to improve the numerical accuracy of the algorithm by reducing the amount of time that the MPOs need to be evolved for.
Overall, the runtime for calculating the OTOC with this method scales as
$\mathcal{O}\left(n\chi^{3}t/\Delta t+n\chi^{4}\right),$ (32)
where $n$ is the size of the quantum system and $t$ is the evolution time. The first term in (32) comes from performing TEBD to evolve the MPOs (see Lemma 2 of [76]), and the second term comes from contracting the MPOs together to calculate the OTOC (see Section III.A.3 in [78]). Compared with the exponential runtime scaling for exact diagonalisation, the runtime for the MPO-based method appears to scale just linearly with the system size $n$. However as the system size increases, we increase the maximum evolution time $t_{max}$ linearly (see Section III.1), and typically a larger $\chi$ is necessary to maintain accuracy, which effectively results in a high degree polynomial runtime scaling in $n$.
Using this method, for each qubit number considered ($n\in\{5,10,15,20,25,30,35,40\}$), each parameterised set of Hamiltonians, and both functions $\mathscr{O}_{XZ}$ and $\mathscr{O}_{Sum}$, we generate 1250 labelled data samples, 1000 for training and 250 for testing. The random samples of the inputs are drawn using the same random seed. Thus, for each qubit number, all datasets contain the same input parameters. The values of the functions are then computed using the ITensor Julia library [79] to implement the MPO-based time-splitting algorithm. It should be noted though, that even with the MPO-based algorithm, producing the data in this work required $\sim 10^{6}$ CPU hours, and thus required access to a supercomputer.